• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Higinbotham

    Log entry time 16:05:17 on February26,2009

    Entry number 263511


    keyword=Shift Summary

    We started the shift with a vacuum problem in the north linac causing the beam studies to start early.

    During the down time, a new reference cell was readied and right Q3 was made operational and Jack worked on the left Q3.

    A quick check of the field shows the sign of the right Q3 is the same as it has always been (checked vs. a 2007 end of run file).